OpenStudy (anonymous):

can a anyone explain how to do this problems? Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

Nnesha (nnesha):

set it equal to 0

Nnesha (nnesha):

let sinx= x write it as quadratic equation \[\huge\rm 7x^2 -14x +2 +5 =0\]

OpenStudy (anonymous):

7sin2x - 14sinx + 2 = -5 7sin2x - 14sinx = -7 7(sin2x - 2sinx) = -7 sin2x - 2sinx = -1 2sinxcosx - 2sinx = -1 2sinx (cosx - 1) = -1 Now solve 2sinx = -1 cosx - 1 ==1

OpenStudy (alekos):

Are we looking at \[\sin ^{2}x\] or sin2x